OpenStudy (anonymous):

Can anyone help me with this one ? find the center, verticles, and foci of this ellipse. 2.(x+6)^2/12+(Y-4)^2/16=1 @ikram002p @dan815 @86camaro @divu.mkr

OpenStudy (ikram002p):

 start wid \[\frac{ x^2 }{ 6 }+\frac{ y^2 }{16 }\]

OpenStudy (ikram002p):

then the center point is (-6,4) so shift that ellips from (0,0 ) to this point.

OpenStudy (anonymous):

the verticles is (+-4sqrt6)

OpenStudy (anonymous):

@ikram002p im confused . Whats the verticles? & whats the foci?

OpenStudy (ikram002p):

 for x^2/6+y^2/16=1

OpenStudy (anonymous):

thanks